A soccer ball is kicked horizontally off a 22m-high hill and lands a distance

of 35.0m from the edge of the hill. Determine the initial horizontal velocity

of the soccer ball.


1
Expert's answer
2022-04-11T14:00:54-0400

The horizontal range of the projectile

"L=v_0\sqrt{\frac{2H}{g}}"

Hence, the initial horizontal velocity

"v_0=L\sqrt{\frac{g}{2H}}=35.0\sqrt{\frac{9.81}{2*22}}=16.5\:\rm m/s"